A turbine operating under steady flow condition receives 5000kg of steam per hour. The steam enter the turbine at a velocity of 3000m/min, an elevation of 5m and a specific enthalpy of 2787kj/kg, it leaves from the turbine at a velocity of 6000m/min an elevation of 1m and a specific enthalpy of 2259kj/kg. Heat losses from the turbine to the surrounding amount to 16736kj/h
Determine the power output of the turbine
1
Expert's answer
2020-06-25T09:35:01-0400

Solution.

"m=5000kg\/h=1.4kg\/s;"

"C_1=3000m\/min=50m\/s;"

"Z_1=5m;"

"h_1=2787kJ/kg=2.787\sdot10^6J/kg;"

"C_2=6000m\/min=100m\/s;"

"Z_2=1m;"

"h_2= 2259kJ/kg=2.259\sdot10^6J/kg;"  

"Q=-16736kJ/h=-4.649\sdot10^3J/s;"

"m(h_1+\dfrac{C_1^2}{2}+Z_1g)+Q=m(h_2+\dfrac{C_2^2}{2}+Z_2g)+W;"

"Q-W=m((h_2-h_1)+\dfrac{C_2^2-C_1^2}{2}+(Z_2-Z_1)g);"

"-4.649\sdot10^3-W=1.4((2.259\sdot10^6-2.787\sdot10^6+\dfrac{100^2-50^2}{2}+(1-5)9.81);"

"-4.649\sdot10^3-W=-1.014\sdot10^6;"

"W=101kJ;"

Answer: "W=101kJ."